OpenStudy (anonymous):

5X - 6 =14 -4 20 15 4

OpenStudy (goformit100):

5x - 6 =14 or, 5x = 14 + 6 or, 5x = 20 or, (5x)/5 = 20/5............... { 5 gets cancel out as present in numerator and denominator } or, x = 20/ 5 or, x = 4..........Ans.